OpenStudy (anonymous):

How do you solve: lim x->0+ sin(x)ln(2x)

OpenStudy (anonymous):

if you replace x by 0 you get \[0\times -\infty \] right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so standard l'hopital trick is to rewrite at either \[\frac{\ln(2x)}{\csc(x)}\] or \[\frac{\sin(x)}{\frac{1}{\ln(2x)}}\]

OpenStudy (anonymous):

Then choose the one where you get 0/0 or infinity/infinity

OpenStudy (anonymous):

that's really helpful! thank you, i know what to do from there

OpenStudy (anonymous):

i think you have to use l'hopital twice. at least it did.

OpenStudy (anonymous):

yeah, I did. also, if i have a question that says : \[\ln ((1-9x)^{1/x})= (blank)(\ln(1-9x))\] what am I suppose to do?